OpenStudy (anonymous):

Solve for the variable: (x/4) + 3= 5

OpenStudy (luigi0210):

Multiply everything by 4 to get rid of the fraction and get: \[\LARGE x+12=20\] Can you solve it from here?

OpenStudy (judygreeneyes):

Multiply by 4 on both sides: X + 12 = 20. Then subtract 12 from both sides: X = 8 You can always plug your answer back in to the original equation to double-check your answer.

OpenStudy (anonymous):

-12=5 (p-2)